OpenStudy (chrisneedshelp):

Simplify -4(x - 7) + (x + 14). -3x - 14 -5x - 14 -3x + 42

OpenStudy (anonymous):

=-4(x-7) + (x+14) =-4x+28+x+14 =-4x+x+28+14 =-3x+42 or simplify to =-x+14

OpenStudy (chrisneedshelp):

thank you ecah

OpenStudy (anonymous):

:D yw

OpenStudy (chrisneedshelp):

lol how did you just boost up 18 levels!?

OpenStudy (anonymous):

no idea how .. and i have no idea what 18 means?? im new here.. :/